# Strict Subgoal Execution: Reliable Long-Horizon Planning in Hierarchical Reinforcement Learning

### v1

@@ Abstract
Long-horizon goal-conditioned tasks pose fundamental challenges for reinforcement learning (RL), particularly when goals are distant and rewards are sparse. While hierarchical and graph-based methods offer partial solutions, they often suffer from subgoal infeasibility and inefficient planning. We introduce Strict Subgoal Execution (SSE), a graph-based hierarchical RL framework that enforces single-step subgoal reachability by structurally constraining high-level decision-making. To enhance exploration, SSE employs a decoupled exploration policy that systematically traverses underexplored regions of the goal space. Furthermore, a failure-aware path refinement, which refines graph-based planning by dynamically adjusting edge costs according to observed low-level success rates, thereby improving subgoal reliability. Experimental results across diverse long-horizon benchmarks demonstrate that SSE consistently outperforms existing goal-conditioned RL and hierarchical RL approaches in both efficiency and success rate.
@@ 1 Introduction
Recent advances in reinforcement learning (RL) have achieved impressive success across various domains [1, 2] . In many real-world applications, agents must achieve specific objectives, motivating the development of goal-conditioned RL (GCRL), where agents learn to reach a designated goal state provided by the environment [3, 4, 5]. Unlike conventional RL methods that rely on carefully designed reward functions, GCRL allows agents to directly pursue target states based on goal specifications. However, in sparse-reward and long-horizon environments, goals are often distant, making exploration difficult and hindering effective learning due to a lack of intermediate guidance.
To address this, hierarchical RL (HRL) decomposes the decision process into a high-level policy that selects subgoals and a low-level policy that executes actions to reach them [6, 7]. Since subgoals are typically closer and more attainable than the final goal, this structure facilitates learning in long-horizon settings. Nevertheless, HRL can still fail when the selected subgoals are too difficult for the low-level policy to reach reliably. To mitigate this issue, graph-based HRL methods have been proposed [8, 9, 10, 11, 12, 13]. These approaches construct a graph over the goal space, where nodes represent states or regions and edges denote feasible transitions. Subgoal selection is guided by shortest paths on the graph, improving success rates in complex tasks. However, even graph-based HRL remains limited in performance when the final goal lies far from the current state, as the high-level policy may require too many steps, resulting in unstable training and poor scalability.
While recent approaches have explored graph-based GCRL without hierarchical structures to mitigate the limitations of high-level policies, such methods often lack the flexibility to handle environments involving multiple goals or diverse reward signals, due to the absence of high-level reasoning [14, 15]. To address these limitations, we propose a novel graph-based HRL framework, Strict Subgoal Execution (SSE), which retains the advantages of high-level planning while enhancing robustness and improving long-horizon goal-reaching by enforcing subgoal completion within a single high-level decision step. Our contributions are summarized as follows:
@@
    HRL Framework with Single-Step Subgoal Execution: We introduce a hierarchical framework that guarantees each subgoal is reached within a single high-level decision, removing step limits and reducing the number of high-level actions required for long-horizon tasks.
@@
    Decoupled Exploration for Goal Space Coverage: A dedicated exploration policy is decoupled from the reward-driven high-level policy to guide the agent through underexplored regions of the goal space, improving spatial coverage and training efficiency.
@@
    Failure-Aware Path Refinement: To enhance subgoal reliability, we adjust graph edge costs based on low-level failure statistics, encouraging path planning to avoid unstable transitions and improving the robustness of subgoal execution.
By ensuring that any high-level subgoal in the map is reachable in a single step, our SSE achieves robust performance even when the final goal lies far across the map. The high-level layer enables sequential decision-making in more complex environments with multiple goals or intermediate objectives. Through extensive evaluation on diverse long-horizon, goal-conditioned tasks, our method consistently demonstrates superior success rates and sample efficiency compared to prior GCRL and HRL approaches, validating the effectiveness of the proposed framework.
@@ 2 Preliminaries
2.1 Universal Markov Decision Process and Goal-Conditioned RL
We consider a universal Markov decision process (UMDP) defined as a tuple $(\mathcal{S},\mathcal{G},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{G}$ is the goal space, $\mathcal{A}$ is the action space, $P$ is the transition dynamics, $R$ is the reward function, and $\gamma\in(0,1]$ is the discount factor [3]. At each time step $t$, the agent observes a goal $g\in\mathcal{G}$ and state $s_{t}$, selects an action $a_{t}\sim\pi(\cdot|s_{t},g)$, and receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1},g)$ and next state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. The goal of GCRL is to learn a goal-conditioned policy $\pi$ that maximizes the expected return $\sum_{t=0}^{H}r_{t}$, where $H$ is the episode length. When $\mathcal{S}\neq\mathcal{G}$, we assume the existence of a mapping $\phi$ such that $\phi(s)\in\mathcal{G}$, allowing the agent to infer goal progress from the current state. The goal $g$ may be fixed or randomly sampled per episode; while random goals aid generalization through diversity, the fixed-goal setting is more challenging due to its singular target.
@@ 2.2 HRL Frameworks and Graph-Based Subgoal Planning in GCRL
In goal-conditioned settings, HRL addresses long-horizon challenges by decomposing the policy into a high-level policy $\pi^{h}$ and a low-level policy $\pi^{l}$ [6, 7]. Every $k$ steps, $\pi^{h}(\cdot|s_{t},g)$ selects a subgoal $\tilde{g}_{t}\in\mathcal{G}$, which $\pi^{l}(\cdot|s_{t},\tilde{g}_{t})$ attempts to reach using an auxiliary reward [16, 17, 18]. However, when subgoals are too distant, the low-level policy may fail to reach them within the given horizon, and distance-based penalties can hinder learning, especially under sparse rewards. To mitigate this, graph-based approaches construct a goal-space graph $G=(V,E)$, where $V$ denotes a set of landmark nodes and $E$ consists of edges weighted by the expected effort required to transition between nodes. Landmark nodes are typically selected either by applying farthest point sampling (FPS) to choose states that are maximally distant from one another [12, 19, 20], or by placing virtual goals uniformly from a predefined grid over the goal space [15]. In general, the edge cost is defined as
@@
where $Q^{G}$ estimates the feasibility of reaching $v_{2}$ from $v_{1}$ under $\pi^{l}$, defined as the value function $Q^{l}$ of the low-level policy or a predefined estimator. Full details on $Q^{G}$ are provided in Appendix B.1.
In graph-based HRL, given a subgoal $\tilde{g}_{t}$, the shortest path from the current state’s embedding $\phi(s_{t})$ to $\tilde{g}_{t}$ is computed via Dijkstra’s algorithm [21], producing a sequence of intermediate waypoints $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$, where each $\mathrm{wp}_{i}\in V$. The low-level policy $\pi^{l}(\cdot|s_{t},\mathrm{wp}_{i})$ guides the agent through these waypoints in order: once $\mathrm{wp}_{i}$ is reached, the next target subgoal for $\pi^{l}$ is updated to $\mathrm{wp}_{i+1}$, continuing until $\tilde{g}_{t}$ is reached. The high-level policy is trained on transitions $(s_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k-1}r_{j},s_{t+k})$ in buffer $\mathcal{B}^{h}$, while the low-level policy $\pi^{l}$ is trained using $(s_{t},\mathrm{wp}_{i},a_{t},r^{l}_{t},s_{t+1})$ in buffer $\mathcal{B}^{l}$, where the low-level reward $r^{l}_{t}=-1$ if the agent has not reached $\mathrm{wp}_{i}$, and $r^{l}_{t}=0$ otherwise [12, 19, 20]. In contrast, certain approaches dispense with high-level subgoals and directly train the low-level policy to reach the final goal $g$, under the assumption that intermediate subgoals are unnecessary [14, 15].
@@ 3 Related work
Advances in Goal-Conditioned RL
GCRL refers to RL settings where the agent is explicitly conditioned on a goal [22, 23, 24]. Modern GCRL methods typically employ Universal Value Function Approximators (UVFA) [3], which facilitate generalization across different goals by learning goal-conditioned policies. A key challenge in GCRL is low data efficiency, as transitions collected under one goal often become unusable for others [25]. To address this, many approaches perform goal relabeling in the replay buffer, transforming failed experiences into successful ones for alternative goals [26, 27, 28, 29, 30]. Beyond random goal sampling, recent work emphasizes structured exploration of the goal space to further improve sample efficiency and avoid local optima. This includes methods that prioritize rare or novel states [31, 32], generate goals using curriculum learning [29, 33], or discover diverse and semantically meaningful goals through unsupervised learning [34, 35, 36].
Subgoal Structuring and Planning in Hierarchical RL
Hierarchical frameworks for GCRL introduces multi-level policy structures to address long-horizon tasks by decomposing them into temporally abstracted subgoals. Foundational work has explored option-based [6] and manager-worker [7] frameworks, demonstrating the effectiveness of hierarchical decomposition in complex decision-making [37, 38, 9, 39]. To improve training stability in off-policy settings, data relabeling techniques have been proposed [40], while subgoal reachability has been enhanced through adjacency-based penalties and adversarial training [4, 16]. Recent advances incorporate graph structures to represent the state space and improve exploration in sparse-reward environments [8, 9, 10, 11, 12, 13]. Other methods leverage graph-based planning to better align high-level decisions with low-level policy execution, mitigating mismatches in temporal abstraction [19].
Structured Graph-Based Approaches for Goal-Conditioned RL
Graph-based approaches in GCRL offer a structured representation of the goal space, enabling planning over discrete landmarks and improved navigation in sparse-reward environments [10, 11, 12]. This direction has evolved through integration with latent modeling [13], hybrid control [41], policy-driven graph construction [42], and hierarchical planning [19]. To improve exploration, recent methods propose frontier-based expansion [20], curriculum strategies for reachability-aware goal selection [14], and virtual subgoal generation for broader coverage [15], all contributing to more efficient learning under sparse rewards.
@@ 4 Methods
4.1 Motivation: Rethinking Subgoal Execution in Graph-Based Hierarchical RL
Figure 1: Agent trajectory comparison in goal space $\mathcal{G}$: (a) conventional HRL with $k$-step limit, (b) proposed SSE with multiple subgoals, and (c) SSE selecting $g$ as the initial subgoal. In (a), reachable range is limited by $k$ steps and the agent continues without fully reaching the subgoal. In contrast, SSE allows arbitrary steps but enforces strict subgoal completion. States are shown as $\phi(s_{t})\in\mathcal{G}$.
Graph-based HRL shows strong potential for long-horizon tasks, yet conventional subgoal execution faces key limitations, particularly in sparse-reward and fixed-goal settings. In this section, we revisit standard designs and motivate a new framework that addresses these issues while preserving the hierarchical structure. Fig. 1(a) illustrates how conventional graph-based HRL operates in a fixed-goal setting. When the graph $G$ contains well-distributed landmark nodes, the high-level policy selects a subgoal $\tilde{g}_{t}$ from the current state $s_{t}$, and the low-level policy follows a waypoint path to reach it within a fixed $k$-step horizon. If the subgoal lies beyond this range, the agent may fail and proceed to the next subgoal $\tilde{g}_{t+k}$, causing variation in $\phi(s_{t+k})$ and destabilizing high-level learning. To mitigate this, some prior methods introduce distance-based penalties to favor closer subgoals, but this design still presents several limitations. In sparse-reward, long-horizon tasks, reachability penalties can dominate the learning signal, hindering high-level optimization. Failures such as collisions or instability often yield inconsistent representations $\phi(s_{t+k})$, degrading high-level updates. Moreover, due to the limited range of the low-level policy, reaching distant goals typically requires many intermediate subgoals, leading to inefficient learning when rewards are infrequent.
To address these challenges, we propose the Strict Subgoal Execution (SSE) framework, which enforces that each subgoal be reached within a single high-level step. This removes fixed step limits and allows subgoals to be placed arbitrarily far from the current state. Fig. 1(b) illustrates this mechanism: at each high-level decision step $t_{i}$ (with $t_{0}=0$), the agent is assigned a subgoal $\tilde{g}_{t}$, and the episode continues only if it successfully reaches $\tilde{g}_{t}$ from $s_{t}$; otherwise, it terminates immediately with no reward. This strict rule fundamentally departs from prior approaches by eliminating transitions resulting from failed subgoal attempts. In doing so, SSE provides more consistent learning signals and stabilizes high-level training. By removing artificial constraints such as step horizons and distance-based penalties, SSE emphasizes reliable subgoal completion as a key learning signal. This leads to more coherent planning in complex, long-horizon environments.
To ensure robust subgoal execution, SSE incorporates two additional components. First, a decoupled exploration policy independently explores under-visited regions and enhances the low-level policy’s ability to reach diverse subgoals, including those that are previously unexplored. Second, a failure-aware path refinement strategy updates edge costs based on subgoal failure ratio rather than geometric distance, promoting more reliable planning. Fig. 1(c) shows the resulting behavior after training with SSE: starting from the initial state, the agent selects the final goal $g$ directly as the first subgoal $\tilde{g}_{0}=g$ and reaches it without relying on intermediate waypoints. This example highlights the expressiveness and reliability of SSE. The proposed framework also generalizes well to complex tasks involving multiple goals or intermediate interactions, often completing them in only a few high-level steps.
4.2 Strict Subgoal Execution with Decoupled Exploration
In this section, we propose the Strict Subgoal Execution framework for graph-based HRL to improve subgoal reachability, as discussed in Section 4.1. While adopting the same graph construction and low-level learning strategies as prior graph-based HRL methods introduced in Section 2, our approach introduces a new high-level policy that ensures subgoal completion without relying on fixed-step limits. In prior works, subgoals are typically selected using Gaussian policies with small noise, often combined with penalties to encourage reachable subgoals. Although suitable under step-limited low-level control, these policies tend to overemphasize nearby goals and thus hinder effective exploration across the goal space. In contrast, our objective is to learn a high-level policy capable of reaching any point on the map in a single decision. To this end, we define an $\epsilon$-greedy high-level policy $\pi^{h}$ as:

### v2

@@ Abstract
Long-horizon goal-conditioned tasks pose fundamental challenges for reinforcement learning (RL), particularly when goals are distant and rewards are sparse. While hierarchical and graph-based methods offer partial solutions, their reliance on conventional hindsight relabeling often fails to correct subgoal infeasibility, leading to inefficient high-level planning. To address this, we propose Strict Subgoal Execution (SSE), a graph-based hierarchical RL framework that integrates Frontier Experience Replay (FER) to separate unreachable from admissible subgoals and streamline high-level decision making. FER delineates the reachability frontier using failure and partial-success transitions, which identifies unreliable subgoals, increases subgoal reliability, and reduces unnecessary high-level decisions. Additionally, SSE employs a decoupled exploration policy to cover underexplored regions of the goal space and a path refinement that adjusts edge costs using observed low-level failures. Experimental results across diverse long-horizon benchmarks show that SSE consistently outperforms existing goal-conditioned and hierarchical RL methods in both efficiency and success rate. Our code is available at https://github.com/Jaebak1996/SSE
@@ 1 Introduction
Recent advances in reinforcement learning (RL) have achieved impressive success across various domains (Mnih et al., 2013; Silver et al., 2016). Although prior works have enhanced sample efficiency through advanced exploration (Burda et al., 2018; Han and Sung, 2021b; a; Jo et al., 2024) and experience replay mechanisms (Schaul et al., 2015b; Han and Sung, 2017), they often struggle to adapt to changing objectives without extensive retraining. Consequently, Goal-Conditioned RL (GCRL) was introduced to explicitly condition the policy on a target goal, allowing agents to directly pursue diverse states specified by the environment (Schaul et al., 2015a; Levy et al., 2017; Nasiriany et al., 2019). However, in sparse-reward and long-horizon environments, goals are often distant, making exploration difficult and hindering effective learning due to a lack of intermediate guidance.
Hierarchical reinforcement learning (HRL) addresses long-horizon goal-conditioned tasks by decomposing control into a high-level policy that proposes subgoals or skills and a low-level policy that executes actions to reach them (Bacon et al., 2017; Vezhnevets et al., 2017). While this decomposition can ease learning by introducing more attainable intermediate objectives, performance can degrade when proposed subgoals are not reliably reachable by the low-level policy. To mitigate this issue, graph-based HRL methods construct a graph over the goal space where nodes represent states or regions and edges encode feasible transitions (Zhang et al., 2018; Nachum et al., 2018a; Huang et al., 2019; Eysenbach et al., 2019; Kim et al., 2021; Zhang et al., 2021). This structure enables subgoal selection via shortest-path planning. Despite these advances, scalability remains challenging when the final goal is far from the current state because high-level planning may require many decisions and lead to unstable training. To reduce reliance on high-level policies, recent work has explored graph-based GCRL without explicit hierarchical structures (Lee et al., 2023; Yoon et al., 2024). However, removing high-level reasoning can reduce flexibility in environments with multiple goals or heterogeneous reward signals.
To overcome these limitations, we propose Strict Subgoal Execution (SSE), a graph-based HRL framework that improves long-horizon performance by explicitly characterizing the reachability boundary of candidate subgoals. SSE integrates Frontier Experience Replay (FER) to distinguish high-level outcomes into failure transitions and partial-success transitions, localizing where execution breaks down and how far it progresses toward a subgoal. This frontier-based separation enables the agent to identify unreliable subgoals, train policies to execute admissible subgoals more accurately, and reduce unnecessary high-level decisions. SSE also expands goal-space coverage by actively exploring undervisited graph regions and improves planning reliability by refining paths to avoid transitions with frequent failures. As a result, SSE solves a broader range of complex long-horizon tasks. Our contributions are summarized as follows:
@@
    Frontier Experience Replay (FER) for SSE: We introduce FER, which delineates the reachability frontier using two high-level samples: failure transitions and partial-success transitions. By precisely localizing where attempts fail and how far progress extends, FER identifies unreliable subgoals, increases subgoal reliability and reduces unnecessary high-level decisions.
@@
    Decoupled Exploration for Goal Space Coverage: A dedicated exploration policy is decoupled from the return-driven high-level policy to traverse underexplored regions of the goal space, improving coverage and sample efficiency.
@@
    Failure-Aware Path Refinement: To improve subgoal reliability, we adjust graph edge costs based on low-level failure statistics, encouraging path planning to avoid unstable transitions and strengthening subgoal execution.
Through extensive evaluation on diverse long-horizon goal-conditioned tasks, our method demonstrates higher success rates and better sample efficiency than prior GCRL and HRL approaches, validating the effectiveness of the proposed SSE framework.
@@ 2 Preliminaries
2.1 Universal MDP, Goal-Conditioned RL, and Goal Relabeling Techniques
We consider a universal Markov decision process (UMDP) $(\mathcal{S},\mathcal{G},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$, $\mathcal{G}$, and $\mathcal{A}$ denote the state, goal, and action spaces, $P$ the transition dynamics, $R$ the reward function, and $\gamma\in(0,1]$ the discount factor Schaul et al. (2015a). At time step $t$, the agent observes $(s_{t},g)$, samples $a_{t}\sim\pi(\cdot|s_{t},g)$, and receives $r_{t}=R(s_{t},a_{t},s_{t+1},g)$ with $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. The goal of GCRL is to learn a goal-conditioned policy $\pi$ that maximizes the expected return $\sum_{t=0}^{H}r_{t}$ over horizon $H$. When $\mathcal{S}\neq\mathcal{G}$, GCRL typically assumes a mapping $\phi:\mathcal{S}\rightarrow\mathcal{G}$ projecting states into the goal space. The goal $g$ may be fixed or sampled per episode. In long-horizon goal-conditioned settings, learning is often inefficient due to sparse positive signals. Goal relabeling methods such as Hindsight Experience Replay (HER) (Andrychowicz et al., 2017) improve sample efficiency by treating later achieved states as substitute goals. HER replaces the goal $g$ with a future achieved goal $g^{\prime}=\phi(s_{t^{\prime}})$ for some $t^{\prime}\geq t$ and recomputes the reward:
$(s_{t},a_{t},r_{t},s_{t+1},g)\ \mapsto\ (s_{t},a_{t},R(s_{t},a_{t},s_{t+1},g^{\prime}),s_{t+1},g^{\prime}).$
This converts unsuccessful attempts into useful signals while the original transitions unchanged.
@@ 2.2 HRL Frameworks and Graph-Based Subgoal Planning in GCRL
In goal-conditioned settings, HRL addresses long-horizon challenges by decomposing control into a high-level policy $\pi^{h}$ and a low-level policy $\pi^{l}$ Bacon et al. (2017); Vezhnevets et al. (2017). In general, HRL frameworks in the GCRL domain typically assume that the goal space $\mathcal{G}$ is known to facilitate the high-level selection of feasible subgoals. Every $k$ steps, $\pi^{h}(\cdot\mid s_{t},g)$ selects a subgoal $\tilde{g}_{t}\in\mathcal{G}$, which $\pi^{l}(\cdot\mid s_{t},\tilde{g}_{t})$ attempts to reach using an auxiliary reward  (Zhang et al., 2020; Pateria et al., 2021; Hutsebaut-Buysse et al., 2022). However, when subgoals are too distant, the low-level policy may fail to reach them within the given horizon, and distance-based penalties can hinder learning under sparse rewards. To mitigate this, graph-based approaches construct a goal-space graph $G=(V,E)$, where $V$ is a set of landmark nodes and $E$ contains edges weighted by the effort to transition between nodes. Landmarks are commonly chosen via farthest point sampling (FPS) (Kim et al., 2021; Lee et al., 2022) or placed on a predefined grid over the goal space (Yoon et al., 2024), assuming that the goal space is known to define the graph and the high-level policy.
In general, the edge cost is defined as
@@
where $Q^{G}$ estimates the feasibility of reaching $v_{2}$ from $v_{1}$ under $\pi^{l}$, defined as the value function $Q^{l}$ of the low-level policy or a predefined estimator. In this paper, we set modified version of edge cost and details on $Q^{G}$ are provided in Appendix B.1. Given a subgoal $\tilde{g}_{t}$, graph-based HRL computes the shortest path from $\phi(s_{t})$ to $\tilde{g}_{t}$ via Dijkstra’s algorithm (Dijkstra, 1959), yielding waypoints $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$ with $\mathrm{wp}_{i}\in V$. The low-level policy sequentially targets these waypoints, switching to $\mathrm{wp}_{i+1}$ once $\mathrm{wp}_{i}$ is reached, until $\tilde{g}_{t}$ is attained. The high-level policy is trained on transitions $(s_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k-1}r_{j},s_{t+k})$ in $\mathcal{B}_{F}^{h}$, while $\pi^{l}$ is trained on $(s_{t},\mathrm{wp}_{i},a_{t},r^{l}_{t},s_{t+1})$ in $\mathcal{B}^{l}$. Following prior work (Kim et al., 2021; Lee et al., 2022), the low-level reward is $r^{l}_{t}=0$ if $\phi(s_{t+1})-\mathrm{wp}_{i}||<\lambda$ and 1-1 otherwise, where $\lambda>0$ is a subgoal reachability threshold and $\phi$ is known. In contrast, some approaches omit high-level subgoals and directly train the low-level policy to reach the final goal $g$, assuming intermediate subgoals are unnecessary (Lee et al., 2023; Yoon et al., 2024).
@@ 3 Related work
Goal-Conditioned RL and Hierarchical Approaches
GCRL refers to RL settings where the agent is explicitly conditioned on a goal (Kaelbling, 1993; Liu et al., 2022; Colas et al., 2022). Modern GCRL typically employs Universal Value Function Approximators (UVFA) (Schaul et al., 2015a) to generalize across goals. A central challenge is solving long-horizon tasks with sparse rewards, where exploration is difficult. To address this, HRL introduces multi-level policies that decompose complex tasks into temporally abstract subgoals (Vezhnevets et al., 2017; Nachum et al., 2018b) or reusable skills (Eysenbach et al., 2018; Lee et al., 2025), thereby reducing the effective planning horizon. The effectiveness of this decomposition has been demonstrated in diverse settings (Barto and Mahadevan, 2003; Kulkarni et al., 2016; Nachum et al., 2018a; 2019). Beyond hierarchical structures, other methods improve sample efficiency through structured exploration, such as prioritizing novel states (Warde-Farley et al., 2018; Pong et al., 2019) or discovering useful goals via unsupervised learning (Mendonca et al., 2021; Ecoffet et al., 2021; Chane-Sane et al., 2021).
Relabeling and Guidance Techniques in GCRL
In sparse-reward GCRL, data relabeling is central for improving sample efficiency. HER converts failed trajectories into successes by replacing the intended goal with a future achieved goal (Andrychowicz et al., 2017). Since uniform hindsight-goal sampling can be suboptimal, subsequent work explores curriculum-based relabeling (Fang et al., 2019), novelty- or priority-driven goal sampling (Zeng et al., 2023), and hierarchical relabeling that better matches low-level behavior (Nachum et al., 2018b). Orthogonal guidance approaches discourage infeasible subgoals, for example via learned adjacency constraints (Zhang et al., 2020) or adversarially generated goals that are challenging yet achievable (Levy et al., 2017).
Graph-based Approaches in GCRL
Graph-based approaches have been introduced to GCRL to provide a structured representation of the goal space, enabling planning over discrete landmarks and improving navigation in sparse-reward environments (Huang et al., 2019; Eysenbach et al., 2019; Kim et al., 2021). Early work used graph structures to represent the state space and to guide exploration (Zhang et al., 2018; Nachum et al., 2018a), and this direction has since evolved through integration with latent modeling (Zhang et al., 2021) and policy-driven graph construction (Kim et al., 2023). More recent advances focus either on aligning high-level decisions with low-level execution via graph-based planning (Lee et al., 2022) or on enhancing exploration with strategies such as frontier-based expansion (Hoang et al., 2021; Park et al., 2024), curriculum-based goal selection (Lee et al., 2023), and virtual subgoal generation for broader coverage (Yoon et al., 2024). In related works, frontier-based exploration refers to exploring near boundary states. In our work, we also use the term frontier, but it denotes the boundary that identifies failure points and separates achievable from unachievable subgoals, which is a distinct objective from frontier-based exploration.
@@ 4 Methods
4.1 Motivation: Rethinking Subgoal Execution in Graph-Based HRL
Figure 1: Agent trajectories in goal space $\mathcal{G}$. (a) Conventional HRL with HER relabels intermediate states as subgoals without enforcing exact subgoal completion, which lengthens high-level trajectories. (b) SSE with FER enforces exact subgoal completion, increasing subgoal reliability and reducing unnecessary high-level decisions, thereby improving learning efficiency. (c) After training, SSE reaches $g$ with few high-level steps, here in a single step in single-goal settings even from distant starts. Agent locations are $\phi(s_{t})\in\mathcal{G}$ and $t_{i}$ is the $i$-th high-level step.
In GCRL, many methods improve goal-reaching performance by adopting hindsight relabeling techniques such as HER (Andrychowicz et al., 2017), which treat intermediate states as virtual goals to provide additional training signals. Recent graph-based HRL methods typically apply this idea to both the low level and the high level. While this benefits the low level, it introduces a critical issue at the high level. Fig. 1(a) illustrates how graph-based HRL with HER operates: the high-level policy selects a subgoal $\tilde{g}_{t}$ given the current state $s_{t}$, and the low-level policy follows a waypoint path to reach it. If $\tilde{g}_{t}$ is unreachable due to limited skill, obstacles, or excessive distance, the agent stops at an intermediate state, and such subgoals should be avoided. When HER is applied, every visited state along the failed trajectory is treated as if it were the intended subgoal. This causes the high-level policy to repeatedly select ineffective subgoals and to waste many decision steps. In addition, the resulting transitions can vary widely for the same subgoal, which hinders stable learning. As shown in Fig. 1(a), this produces unnecessarily long high-level trajectories, and even if the agent reaches the goal $g$, credit is spread over too many steps, preventing earlier decisions from being reinforced and often leading to failure on long-horizon tasks.
To address this issue, we propose the Strict Subgoal Execution (SSE) framework, which updates the high-level policy with positive returns only when the low level successfully reaches the assigned subgoal. We instantiate this principle with Frontier Experience Replay (FER). Unlike relabeling methods such as HER that synthesize additional successes, FER delineates the reachability frontier by recording two types of high-level samples, failure transitions and partial-success transitions. By precisely localizing where attempts fail and how far progress extends, FER identifies unreliable subgoals, provides consistent training signals, and reduces unnecessary high-level decisions. As shown in Fig. 1(b), this separation of success and failure keeps the resulting state $\phi(s_{t^{\prime}})$ closely aligned with the selected subgoal $\tilde{g}_{t}$, producing consistent high-level transitions and eliminating wasteful actions. Consequently, as illustrated in Fig. 1(c), the resulting high-level policy solves tasks with far fewer decisions, often reaching the final goal $g$ in a single step in simple settings and handling multi-goal or long-horizon environments with only a few well-chosen subgoal selections.
4.2 Strict Subgoal Execution with Frontier Experience Replay
In this section, we describe the details of the SSE framework. We first define FER, the key component of SSE, which marks the reachability frontier by recording two high-level sample types in addition to standard successes: failure transitions that stop at the point of failure and partial-success transitions that record the last reliably reached waypoint. To formalize this, we basically follow the HRL setup from Section 2: At each time $t$, the policy selects a subgoal $\tilde{g}_{t}\in\mathcal{G}$ under the common assumption that the goal space $\mathcal{G}$ is known. A waypoint path $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$ on the graph $G=(V,E)$ is then generated, and the low-level policy $\pi^{l}$ follows this path until termination at time $t^{\prime}$. We regard the subgoal as reachable if $\|\phi(s_{t^{\prime}})-\tilde{g}_{t}\|<\lambda$; otherwise the attempt is treated as a failure. In the failure case, $\mathrm{wp}_{\mathrm{final}}$ denotes the last waypoint reached within tolerance before failure. Based on this notion of reachability, we define FER as follows.
Definition 4.1 (Frontier Experience Replay)
The high-level replay buffer $\mathcal{B}_{F}^{h}$ stores transitions as
[…]
Here, $s_{T}$ is the terminal state, $\sum_{j}r_{j}$ is the cumulative reward collected until reaching the subgoal or waypoint, and $t_{\mathrm{wp}}$ is the time step at which $\mathrm{wp}_{\mathrm{final}}$ is reached.
In FER, a success transition records the full cumulative return up to $t^{\prime}$ and continues the episode, a stop-on-failure transition assigns zero return and sets the next state to $s_{T}$, which immediately truncates the episode so all future returns are forfeited and unreachable or unreliable subgoals are discouraged, and a partial-success transition records $\mathrm{wp}_{\mathrm{final}}$ with its accumulated return, localizing how far the attempt progressed and crediting only the reliably executed portion. Together, these signals delineate the reachability frontier, identify unreliable subgoals, and provide consistent high-level targets that reduce unnecessary high-level decisions. This setup separates reliable from unreliable subgoals and suppresses failure-inducing actions, but it can induce conservatism and reduce exploratory coverage because the agent learns to avoid regions associated with failure. To counterbalance this effect, we introduce a decoupled high-level controller comprising a high-level policy $\pi^{h}$ for exploitation trained on $\mathcal{B}_{F}^{h}$ and a complementary exploration policy $\pi^{\mathrm{exp}}$ that promotes coverage. As in prior graph-based HRL, we assume the goal space $\mathcal{G}$ is known for subgoal selection of both $\pi^{h}$ and $\pi^{\mathrm{exp}}$, and SSE introduces no additional assumptions. Detailed formulations follow.
High-level Policy $\pi^{h}$: In prior work, subgoals are typically selected with Gaussian policies with small noise. Although adequate for those methods, this concentrates exploration near the current maximum subgoal. As noted above, we aim for a high-level policy that can target any point in the goal space, thereby broadening exploration. We therefore define an $\epsilon$-greedy $\pi^{h}$ as